# What Evidence Do Language Models Find Convincing?

### v1

@@ Appendix A Additional Details on ConflictingQA
Table 5 lists each question category in the dataset and Table 6 contains the prompt used to generate these category. Table 7 contains the prompt used to classify the stance of the retrieved websites.
Publishing, Biodiversity, Religion, Endangered Species, Pomology, Odontology, Pharmacology, Diabetology, Lepidopterology, Horticulture, Paleoclimatology, Product Design, Sustainability, Genomics, Intellectual Property, Gemology, Biomathematics, Karyology, Biomechanics, Selenology, Meteoritics, Chronobiology, Online Learning, Sustainable Living, Mammalogy, Web Design, Cytogenetics, Politics, Veterinary Science, Informatics, Zoogeography, Organic Farming, Cryptocurrency, Ethnobotany, Petrology, Serology, Ethology, Seismology, Entrepreneurship, Zymology, Astronomy, Holistic Health, Ichthyology, Trichology, Hematology, Gerontology, Neurology, Aging, Heuristics, Nematology, Nuclear Energy, Conservation, Botany, Spelaeology, Urology, Virology, Ergonomics, Volcanology, Yoga, Dermatology, Stomatology, Paleopathology, Xenobiology, Anthropometry, Anthropology, Entertainment, Poetry, Animation, Archaeology, Ornithology, Radio, Etymology, World Religions, Oncology, Anthrozoology, Criminology, Herpetology, Television, Malacology, Paranormal, Philology, Forestry, Probabilistics, Aerospace, Somnology, Cardiology, Cognitive Science, Quantum Physics, Phylogenetics, Epistemology, Vulcanology, Epidemiology, Psychobiology, Kinematics, Telecommunications, Melittology, Otorhinolaryngology, Astronautics, Biophysics, Neuroscience, Paleo Diet, Endocrinology, Kinesiology, Constitutional Law, Pop Culture, Lexicology, Festivals, Evolution, Metallurgy, Pediatrics, Phonetics, Astrobiology, Pets, Multiculturalism, Veganism, Andragogy, Remote Work, Speleology, Telepathy, Marine Conservation, Human Geography, Creationism, Philosophy, Oceanography, Mycology, Work-Life Balance, Ethics, Bioethics, Viniculture, Pedagogy, Classical Music, Paleoethnobotany, Manuscripts, Paleobotany, Revolutions, Paleozoology
Table 5: The full list of the 136 categories from Table 1.
@@
Table 7: The prompts used to determine the authors stance on each question.
@@ Appendix B Model-based Data Filtering
We filter out any paragraph that the downstream LLM predicts a different stance for than the ensemble of GPT-4 and Claude v1 Instant. We do this by taking the paragraph of interest and comparing it to a paragraph with the text “This website has no text”. We remove any paragraph where the model’s output differs from the predicted stance label. We also balance the dataset such that each answer (Yes or No) to a question contains an equal number of convincing and unconvincing paragraphs.
@@ Appendix D Expressing Convincingness in Isolation
We consider whether LLMs are able to express the convincingness of a paragraph in isolation. The model makes the rating using only the website. We prompt (Table 9) asking the model to rate the credibility of the website from a scale of one to five. The rating of the model is then determined by an average of the ratings, weighted by the probability of each label. Following Santurkar et al. (2023), we calculate probabilities by exponentiating and normalizing the logits for "one" through "five". We also give the model with examples of a "one" and "five" ratings from C3 Kakol et al. (2017), a dataset for studying human credibility judgements. We use these few-shot examples as the model tended to be biased toward higher-ratings without them.

### v2

@@ Appendix A Additional Details on ConflictingQA
Table 5 lists each question category in the dataset, and Table 6 contains the prompt used to generate these categories. Table 8 contains the prompt used to convert questions to affirmative and negative statements. Table 9 contains the prompt used to classify the stance of the retrieved websites.
Publishing, Biodiversity, Religion, Digital Rights, Endangered Species, Biotechnology, Pomology, Virtual Reality, Numismatics, Wilderness Exploration, Entomology, Pharmacology, Diabetology, Ornithology, Lepidopterology, Horticulture, Ethology, Paleoclimatology, Product Design, Seismology, Climate Change, Sustainability, Stomatology, Rhetoric, Genomics, Intellectual Property, Gemology, Biomathematics, Philosophy, Karyology, Biomechanics, Telecommunications, Selenology, Meteoritics, Demographics, Chronobiology, Malacology, Marine Conservation, Online Learning, Agribusiness, Sustainable Living, Ecophysiology, Mammalogy, Herpetology, Politics, Web Design, Cytogenetics, Neuroscience, Bioacoustics, Veterinary Science, Informatics, Zoogeography, Organic Farming, Cryptocurrency, Ethnobotany, Data Privacy, Petrology, Real Estate, Rheumatoid, Serology, Epistemology, Astronomy, Entrepreneurship, Zymology, Melittology, Pets, Probabilistics, Holistic Health, Evolution, Ichthyology, Aging, Trichology, Hematology, Gerontology, Hydrology, Neurology, Metallurgy, Heuristics, Nematology, Nuclear Energy, Conservation, Botany, Dermatology, Renewable Energy, Robotics, Spelaeology, Gastroenterology, Psychobiology, Urology, Creationism, Paleo Diet, Virology, Ergonomics, Veganism, Volcanology, Folklore, Yoga, Paleopathology, Speculative Fiction, Xenobiology, Anthropology, Theater, Paleobotany, World Religions, Pop Culture, Anthropometry, Entertainment, Ancient Civilizations, Poetry, Comics, Animation, Festivals, Archaeology, Dance, Radio, Etymology, Sports, Otorhinolaryngology, Mycology, Oncology, Anthrozoology, Criminology, Television, Paranormal, Philology, Forestry, Aerospace, Somnology, Broadcasting, Cardiology, Cognitive Science, Quantum Physics, Phylogenetics, Vulcanology, Epidemiology, Nephrology, Kinematics, Astronautics, Biophysics, Endocrinology, Kinesiology, Odontology, Pediatrics, Vaccinology, Semiotics, Thermodynamics, Constitutional Law, Viniculture, Metaphysics, Lexicology, Astrobiology, Civil Rights, Plastic Surgery, Typography, Venereology, Networking, Cryptanalysis, Advertising, Graphic Design, Cloud Computing, Dacryology, Data Science, Thanatology, Toxicology, Human Geography, Transportation, Etiquette, Public Transport, Phonetics, Neuropathology, Multiculturalism, Andragogy, Remote Work, Speleology, Telepathy, Algorithms, Sociology, Bibliography, Oceanography, Work-Life Balance, Ethics, Bioethics, Endoscopy, Pedagogy, Cartography, Classical Music, Paleoethnobotany, Manuscripts, Ufology, Revolutions, Paleozoology
Table 5: The full list of the 191 categories from Table 1.
@@
Question: {question}
This is a yes-or-no question. Rewrite this question as a statement, first in the affirmative and then in the negative.
Write multiple versions of each statement, varying the wording and phrasing.
Use the following format for your answer:
Yes 1: Statement where answer is Yes
Yes 2: Another statement where answer is Yes
No 1: Statement where answer is No
No 2: Another statement where answer is No
Example input:
Question: Is red wine good for the heart?
Example answer:
Yes: Red Wine Is Good for the Heart
No: Red Wine Is Not Good for the Heart
Table 8: The prompt used to convert questions into statements. We generate multiple rephrasings, then concatenate them together to form the affirmative and negative statements.
@@
Table 9: The prompts used to determine the author’s stance on each question.
@@ Appendix B Model-based Data Filtering
We filter out any paragraph that the downstream LLM predicts a different stance for when compared against the ensemble of GPT-4 and Claude v1 Instant. We do this by making a paired prediction (Table 10) using the paragraph of interest and the text “This website has no text”. We remove any paragraph where the model’s output differs from the stance label.
@@ Appendix D Expressing Convincingness in Isolation
We consider whether LLMs are able to express the convincingness of a paragraph in isolation. The model makes the rating using only the website. We prompt (Table 7) asking the model to rate the credibility of the website from a scale of one to five. The rating of the model is then determined by an average of the ratings, weighted by the probability of each label. Following Santurkar et al. (2023), we calculate probabilities by exponentiating and normalizing the logits for "one" through "five". We also give the model examples of a "one" and "five" rating from C3 Kakol et al. (2017), a dataset for studying human credibility judgements. We use these few-shot examples as the model tends to be biased toward higher-ratings without them.
@@ Appendix E Counterfactual Perturbations
We include the prompts used to make these perturbations in Table 11. We additionally measure the maximum n-gram overlap between the query and the perturbed samples to confirm that the altered paragraphs are, in fact, more relevant to the query. These results can be found in Table 12.
@@
Perturbation Prompt
Rewrite Tech. Language Please alter this text so that it uses technical language. Do not alter the underlying content of the text. Only alter the style of the text. Answer with nothing but the changed text. Only answer with the altered text.
Add More Info Please add 6 sentences to this text that are unrelated to "{question}", but related to the overall topic of the text. The added six sentences should not influence the answer of the "{question}" one way or the other. The added 6 sentences should be related to the overall topic of the text. The added 6 sentences should be sprinkled out throughout the text. Do not put ALL the sentences at the beginning or end. Answer with nothing but the changed text.
Add Science Reference Please alter this text so that it include references to scientific sources. Do not alter the underlying content of the text. Only alter the style of the text. Answer with nothing but the changed text. Only answer with the altered text.
Rewrite Objective Please alter this text so that it sounds more objective. Do not alter the underlying content of the text. Only alter the style of the text. Answer with nothing but the changed text. Only answer with the altered text.
Rewrite Confidence Please alter this text so that it is extremeley confident. Each sentence should be clear and unambiguous. Do not alter the underlying content of the text. Only alter the style of the text. Answer with nothing but the changed text. Only answer with the altered text.
Keyword Stuffing Please alter this text so that it uses keywords related to the question by additional sentences. Do not alter any existing sentences. Existing sentences should remain unaltered. Just add new sentences that use keywords related to the question. Do not alter the underlying content of the text. Only alter the style of the text. Answer with nothing but the changed text.
Add Single Sentence Please minimally alter this text so that the answer to the question "{question}" is obviously {stance}. You are only allowed to add a single sentence. Only add a single sentence and change nothing else. Answer with nothing but the changed text.
Rewrite Relevance Please alter this text so that it is more related to the question "{question}". Emphasize sentences that answer the question "{question}". Add a lot of keywords related to the question into the text. The text should use a lot of keywords related to the question "{question}". Answer with nothing but the changed text.
Table 11: The prompts used to perform counterfactual perturbations on the text.
Perturbation Type Overlap
Rewrite Relevance (Claude) Relevance 5.08
Add Single Sentence (Claude) Relevance 3.64
Question Prefix Relevance 6.89
Keyword Stuffing (Claude) Relevance 2.47
Add Science Reference (Claude) Style 2.38
Add More Info (Claude) Style 2.46
Add Contact Info Style 2.41
Rewrite Confidence (Claude) Style 2.17
Rewrite Tech. Language (Claude) Style 1.97
Rewrite Objective (Claude) Style 2.15
Table 12: We calculate the maximum length n-gram that’s in both the search query and the perturbed paragraph. We confirm that the paragraphs altered by Claude to be more relevant to the search query are also measured to be more relevant.